News & Status
  • News & status not loaded yet.

Javascript is turned off in your web browser.

To take full advantage of this site, please turn it on and then refresh the page.

2017-07-05: New server TempStorage4

The Microscope Facility has a new server for its users: TempStorage4 (>3x bigger & >6x faster). Please delete your shortcut to the old server and create a shortcut to the new one at \\162.129.230.229\TempStorage4 for Windows and //162.129.230.229/TempStorage4 for Macs (see instructions if you've forgotten how). The login and password to the new server is the same as before. Please copy your data from the old server (accessible as a read-only subdirectory 'TempStorage3' on TempStorage4) to the new server.

6/30/2017 11:41am>>
TempStorage3 Offline/to Retire
TempStorage3 is offline while we bring up the new Server (Tempstorage4), which is >3x bigger and >6x faster. You will need to edit your shortcuts to the new folder (Windows: "\\162.129.230.229\Tempstorage4"; Mac: "162.129.230.229/Tempstorage4"), noting that the 's' is lowercase. The login/password is the same as before for TempStorage3, but will no longer be accessible.

If you don't know how to create such shortcuts, see FastFacts/Network.

Direct access to TempStorage3 (SMB1 or FTP) will no longer be available. Your old data will still be accessible through the new server as a linked folder called 'TempStorage3' and you can copy your old data folder from the old server to the new server. Eventually, TempStorage3 will shut down and retire (linked folder will disappear), but you'll have at least one month to transfer your data.


Microscope Facility News & Equipment Status

To change News view:
Two months
Quarter
Year
All
Select multiple with control-key
Restrict listing
to selected
equipment
(Exclude announcements)

2018-01-17 (Leo FESEM): Leo FESEM: Not available until further notice

The Leo FESEM is not available due to chiller issues. As a result, we had to shut the microscope down. We will keep everyone updated and let you know as soon as we have the microscope back up and running.

2018-01-12 (Zeiss LSM 780 FCS): Zeiss LSM780, 561 laser replaced

2018-01-05 (Leo FESEM,Philips CM120 TEM): Chiller swap: LEO offline, Philips working again.

This afternoon, we swapped the chiller for the LEO FESEM with the chiller for the Philips CM120, which is broken. Hence, the Philips is now operational, but the LEO is now offline.

Until the chiller gets repaired/replaced (waiting for technician visits), we plan to time-share the chiller between the LEO and other instruments.

Sorry for the inconvenience.

2018-01-02 (Philips CM120 TEM): Chiller broken

Fri 1/5/18 Update: We swapped the chiller from the LEO SEM to the Philips CM120 so that the Philips is now functional and the LEO offline. The broken chiller was examined by JHU Facilities, but they can't fix it. A contractor will come next week to see what they can do.

Wed 1/2/18: The chiller for the Philips has stopped working. Both the water pump and the condenser are not running, despite being powered. We're waiting for JHU Facilities to examine the equipment and provide their assessment.


Last modified